# Draft Model Knows When to Stop: A Self-Verification Length Policy for Speculative Decoding

### v1

@@ 3.2 Long-form Generation
Most existing works on speculative decoding (Chen et al., 2023; Du et al., 2024) limit their experiments to generating short sequences of 128 tokens. To verify the wide applicability of SVIP, we also conduct experiments on long-form generation with up to 8K context. For this purpose, we use MT-Bench (Zheng et al., 2023) as the dataset and set the sampling temperature to 1, as we found that when using greedy decoding in long-form generation, both the draft and the target models are prone to repeat themselves, resulting in very low information entropy (see Appendix D for details). Other settings follow Section 3.1.
Methods Generation Length
128 256 512 1K 2K 4K 6K 8K
Pythia (6.9B, ${}_{\text{(6.9B, 160M)}}$ Const. 0.68 0.69 0.69 0.70 0.72 0.90 0.89 0.87
Heuristics 0.88 0.88 0.88 0.88 0.90 1.21 1.25 1.23
SVIP 1.07 1.08 1.08 1.07 1.08 1.43 1.44 1.41
Qwen2.5 (14B, ${}_{\text{(14B, 0.5B)}}$ Const. 0.98 0.97 0.95 0.96 0.98 1.00 1.02 1.04
Heuristics 1.01 0.99 0.98 1.00 1.02 1.03 1.04 1.06
SVIP 1.29 1.29 1.30 1.31 1.32 1.33 1.33 1.35
LLaMA-3 (70B, ${}_{\text{(70B, 8B)}}$ Const. 1.74 1.74 1.77 1.78 1.82 1.90 1.93 1.94
Heuristics 1.53 1.56 1.61 1.63 1.68 1.77 1.83 1.84
SVIP 1.69 1.72 1.75 1.78 1.86 1.96 2.01 2.02
Table 2: Speedup on MT-Bench with different generation length.
The results are given in Table 2. Interestingly, we find that in the sampling setting, the two baseline methods (constant and heuristics) perform even worse than target-model-only auto-regressive decoding for Pythia and Qwen, while SVIP consistently yields a positive speedup. For LLaMA, while constant draft length performs better with contexts shorter than 1K tokens, SVIP exceeds it for generating longer sequences. Another observation is that the speedup ratio of all three methods generally increases with the context length, which could be possibly attributed to the longer contexts giving the draft model more information, aligning it at test time to the target model. However, we note that the absolute values of token throughput (which are not discussed in this paper, since they heavily depend on the underlying machines) stay at the same level from 1K to 8K context for Qwen2.5 and LLaMA-3 when using speculative decoding, and decrease slowly when not using speculative decoding. For Pythia, which does not use any optimized attention such as MQA (Shazeer, 2019) or GQA (Shazeer, 2019), the throughput decreases notably with context length in both cases - with or without specualtive decoding.
3.3 Applying SVIP to Other Draft Methods
@@
In Section 3.1 and 3.2, we evaluated SVIP on vanilla speculative decoding, where a standard pretrained Transformer decoder model from the target model’s family is used as the draft model. However, in the past years many works on speculative decoding have proposed other stronger or more efficient draft models (Cai et al., 2024; Du et al., 2024; Li et al., 2024b). Since most of these works assume a constant draft length, SVIP is orthogonal to them and can be applied on top of them without any additional training.
Methods MT-Bench Code Finance GSM Spider Avg.
7B GliDe 1.95 2.04 1.91 1.98 1.69 1.95
   +SVIP 2.00 2.12 2.03 2.01 1.63 2.02
GliDe + CaPE 2.36 2.57 2.29 2.51 1.97 2.40
   +SVIP 2.56 2.65 2.49 2.54 2.08 2.52
13B GliDe 2.22 2.41 2.15 2.31 1.85 2.24
   +SVIP 2.31 2.43 2.17 2.35 1.85 2.28
GliDe + CaPE 2.73 2.86 2.66 2.80 2.24 2.73
   +SVIP 2.72 2.93 2.66 2.85 2.27 2.76
33B GliDe 2.12 2.25 2.09 2.29 1.99 2.18
   +SVIP 2.29 2.40 2.20 2.42 2.03 2.30
GliDe + CaPE 2.08 1.98 2.10 2.13 1.76 2.03
   +SVIP 2.13 2.02 2.15 2.16 1.82 2.08
Table 3: Speedup comparison with GliDe & CaPE, using Vicuna as the base model.
Specifically, we consider GliDe with a CaPE (Du et al., 2024), where the draft model - named GliDe - is a small transformer decoder with cross-attention to the target model’s hidden representations, while CaPE is a complementary tree expansion method to increase the acceptance rate of the draft token at each position. Following the settings of Du et al. (2024), we use Vicuna 7B, 13B, and 33B (Chiang et al., 2023) as the base models (for which the draft models are publicly available) and set the sampling temperature to 0. However, we distribute the 33B model across two GPUs due to memory constraint. Similar to the previous experiments, we set the threshold $t$ to 0.5 based on pilot experiments on 8 samples held out from MT-Bench using GliDe only (without CaPE).
We also apply SVIP to EAGLE-2 (Li et al., 2024a), the state-of-the-art speculative decoding system which utilizes the target model’s language modeling head on top of the draft model’s features to predict the next draft token, and dynamically constructs a draft tree at each draft position. Following Li et al. (2024a), we use LLaMA-2 7B, 13B (Touvron et al., 2023) and Vicuna 7B, 13B (Chiang et al., 2023) as the base models, and set the sampling temperature to 1. As Brown et al. (2024) suggest, adding a conditional clause (that decides whether or not to stop drafting) after the generation of each draft token in EAGLE-2 may interrupt the otherwise input-independent control flow of EALGE-2, reducing the effects of low-level interpreter and system optimizations. Thus, in EAGLE-2 we calculate the draft entropy and decide whether or not to stop drafting after every two draft tokens instead of after every draft token, as used in previous experiments.
Methods MT-Bench H-Eval GSM8K Alpaca CNN/DM QA Avg.
LLaMA-2 7B EAGLE-2 3.10 3.61 3.15 3.10 2.76 2.84 3.10
   + SVIP 3.16 3.66 3.18 3.13 2.82 3.02 3.16
LLaMA-2 13B EAGLE-2 3.38 4.12 3.41 3.25 3.01 2.98 3.41
   + SVIP 3.41 4.09 3.45 3.34 3.05 3.17 3.46
Vicuna 7B EAGLE-2 2.66 2.84 2.77 2.48 2.31 2.13 2.58
   + SVIP 2.84 2.97 2.75 2.66 2.42 2.30 2.73
Vicuna 13B EAGLE-2 2.85 3.31 2.93 2.74 2.48 2.30 2.83
   + SVIP 2.94 3.49 3.19 2.89 2.60 2.53 2.98
@@
Table 4: Speedup comparison with EAGLE-2, using LLaMA-2-Chat and Vicuna as the base models.
The results for these two methods are presented in Table 3 and 4, respectively. For GliDe, SVIP yields a consistent speedup both with and without CaPE, with a 5% improvement for the 7B and 33B models. For EAGLE-2, the speedup is also particularly notable for the Vicuna models, with 5% improvement for both the 7B and 13B models.
@@ 4 Related Work
Works in this category are few, but most relevant to ours. Liu et al. (2024) introduce PEARL, which lets the target model perform verification in parallel to draft generation, stopping the draft process when a mismatch is found. Huang et al. (2024) propose SpecDec++, which trains an acceptance prediction head on top of the draft model to predict the acceptance probability of the current draft token, stopping the draft round when the predicted acceptance probability falls below a constant threshold. Brown et al. (2024) propose Dynamic Depth Decoding on top of EAGLE-2, which uses the sum of all tokens’ confidences in one level of its draft tree as an indicator to predict whether or not to continue draft generation.
@@ 5 Conclusion
We propose SVIP, a flexible, training-free, and plug-and-play dynamic draft length policy for speculative decoding systems. Based on a theoretical lower bound of acceptance probability and its empirical approximation, SVIP determines whether to continue draft generation or to quit drafting based on the draft model’s entropy after the generation of each draft token. With extensive experiments spanning various base models, draft methods, test domains, and generation length, we validated the effectiveness of SVIP, sparking new insights on speculative decoding and more efficient large language models.
@@ Appendix A The Complete Speculative Decoding Algorithms
22:$x_{\leqslant n}$
@@ Appendix B Alternatives for Acceptance Rate Lower Bound Computation
$\geqslant 1-\sqrt{\frac{1}{2}KL(q||p)}.$ (8)
@@
$\beta$ $\geqslant 1-\sqrt{1-e^{-KL(q||p)}}.$ (9)
Compared with the Pinsker’s bound, it’s trivial to see that this bound is guaranteed to be always larger than 0. However, in practice we find that the Pinsker’s bound is 11% tighter for Qwen2.5, 20% tighter for Pythia, and 43% tighter for LLaMA-3.
Appendix C Additional Results on SpecBench
Methods MT-Bench Trans. Sum. QA Math RAG Avg.
[…]
Heuristics 0.82 0.83 0.85 0.83 0.83 0.83 0.83
SVIP 1.05 1.02 1.03 1.01 1.03 1.00 1.02(+56.9%)
[…]
Heuristics 1.02 0.94 0.93 0.88 1.22 0.91 0.99
SVIP 1.24 1.08 1.19 1.11 1.47 1.10 1.20(+25.0%)
[…]
Heuristics 1.56 1.55 1.76 1.49 1.61 1.55 1.58
SVIP 1.53 1.53 1.69 1.51 1.71 1.56 1.58(-1.3%)
Table 5: Speedup on SpecBench using temperature sampling.
In Figure 6, we plot the draft length and acceptance rate of Pythia (which complements Figure 5 in Section 3.1), and in Figure 7 we also give the results of the three models when using sampling instead of greedy decoding [^3] We note that in the “constant” scenario, the draft length is always set to 5. However, if an EOS token is sampled from the draft model, the draft process will terminate. So the overall draft length might be slightly lower than 5, as can be seen from Figure 7.. Across the different models and sampling methods, the same observation as discussed in Section 3 holds: SVIP results in short draft lengths and a much higher acceptance rate.
Figure 6: The average generated draft length, accepted draft length, and acceptance rate of Pythia on SpecBench.
Figure 7: The average generated draft length, accepted draft length, and acceptance rate of Qwen2.5 (top), Pythia (middle), and LLaMA-3 (bottom) on SpecBench, using temperature sampling instead of greedy decoding.
@@ Appendix D Additional Results on Long-form Generation
Methods Generation Length
@@
Pythia (6.9B, ${}_{\text{(6.9B, 160M)}}$ Const. 1.10 1.30 1.50 1.66 1.82 1.20 1.11 1.04
Heuristics 1.25 1.45 1.65 1.81 2.02 1.50 1.46 1.41
SVIP 1.41 1.62 1.83 2.01 2.21 1.71 1.60 1.52
Qwen2.5 (14B, ${}_{\text{(14B, 0.5B)}}$ Const. 1.05 1.08 1.15 1.29 1.44 1.54 1.60 1.67
@@
LLaMA-3 (70B, ${}_{\text{(70B, 8B)}}$ Const. 2.06 2.18 2.31 2.45 2.58 2.72 2.77 2.78
Heuristics 2.26 2.46 2.73 3.07 3.48 3.90 4.15 4.26
SVIP 2.31 2.56 2.86 3.21 3.59 4.00 4.23 4.33
Table 6: Speedup on MT-Bench with different generation length using greedy decoding.
Figure 8: The correlation between draft model entropy and lengths of accepted draft seqeunces in the greedy setting.
In Table 6, we present the results of long-form generation with greedy decoding. Compared with Table 2, the speedup ratio of greedy decoding is much higher (even more than 4 times for LLaMA-3 after 4K context). This is due to the fact that these models tend to repeat themselves in greedy long-form generation, making it very easy for the draft models to predict the next tokens. In Figure 8, we also plot the relation between draft model entropy and accepted draft sequence lengths in this setting, which shows a much stronger correlation compared with the sampling setting in Figure 2.

### v2

@@ 3.1 Results on Long-form Generation
We first validate the effectiveness of SVIP on the widely used MT-Bench (Zheng et al., 2023) using different sizes of Qwen2.5 (Yang et al., 2024) as target and draft models. Unlike many existing works on speculative decoding (Chen et al., 2023; Du et al., 2024) that limit their experiments to generating short sequences of 128 tokens, we conduct experiments on long-form generation with up to 8K context to investigate the applicability of speculative decoding in a broader scope. We set the sampling temperature to 1, as we found that when using greedy decoding in long-form generation, both the draft and the target models are prone to repeat themselves, resulting in very low information entropy and exaggerated speedup ratios (Ouyang et al., 2024) (see Appendix D for details). The entropy threshold $h$ in SVIP is chosen from $\{0.2,0.3,0.4,0.5\}$ based on performance on 8 held out samples using the 14B model as target and the 0.5B model as draft, which is set to 0.3 and reused in all following experiments.
As evaluation metrics, we mainly report the average speedup over target-model-only autoregressive decoding, but also consider other auxiliary information including accepted draft lengths and draft token accept rate. Also, since the memory consumption of verifying $n$ draft tokens is quadratic in $n$, we limit the maximum draft length to 40 in both heuristics and SVIP scenarios, beyond which we start to encounter out-of-memory issues.
SVIP outperforms all baselines
We validate our proposed method, SVIP, with two target models: Qwen2.5-14B and Qwen2.5-32B, utilizing draft models that vary in size from 0.5B to 3B. We show the performance of SVIP and baselines in Figure 5. It shows that SVIP consistently outperforms constant and heuristics draft length by a large margin.
Figure 6: Analysis of Qwen2.5 14B/0.5B’s behaviours on MT-Bench. Compared with constant and heuristics length policies, SVIP generates shorter drafts with a significantly higher accept rate.
Figure 7: Further analysis of Qwen2.5 14B/0.5B on MT-Bench: (a) SVIP outperforms all constant length policies ranging from 2-5; (b) delta draft length (defined as proposed draft length minus oracle draft length) show that constant and heuristics length policies tend to over-generate drafts, while SVIP models the oracle draft length almost perfectly.
To further investigate the origin of SVIP’s speedup, we analyze the proposed draft lengths and accepte rate of the different length policies in Figure 6. We observe that by terminating the draft process early when the draft model entropy is high, SVIP leads to shorter draft lengths and a much higher acceptance rate. However, in Figure 7 we also compare the performance of SVIP with shorter constant draft length policies, which suggest that simply using shorter constant draft length does not suffice, highlighting the importance of dynamically determining draft lengths. Figure 6 also suggests that while the heuristics draft length policy tends to produce very long drafts at long context, its acceptance rate remains low, resulting in no effective speedup compared with fixed draft length.
SVIP  drafts fit oracle length
In Figure 7, we revisit the concept of “oracle draft length” introduced in Section 2.1, and plot the differences between actual draft lengths and oracle draft lengths. The results suggest that both constant and heuristics draft length policies tend to generate drafts that are too long, while SVIP models the oracle draft length almost perfectly, with an average delta below 0.5 tokens.
SVIP further boosts strong baseline
@@
In the previous experiments, we evaluated SVIP on vanilla speculative decoding, where a standard pretrained Transformer decoder model with the same vocabulary as the target is used as the draft model. However, in the past years many works on speculative decoding have proposed other stronger or more efficient draft models (Cai et al., 2024; Du et al., 2024; Li et al., 2024b). Since most of these works assume a constant draft length, SVIP is orthogonal to them and can be applied on top of them without any additional training.
@@
Table 2: Speedup on MT-Bench on top of EAGLE-2, using Vicuna as base models.
Model Method Context Length
128 256 512 1K 2K 4K
V-7B E2 2.70 2.62 2.52 2.41 2.43 1.24
+ SVIP 2.80 2.76 2.71 2.69 2.75 1.41
V-13B E2 2.95 2.90 2.83 2.74 2.71 1.53
+ SVIP 2.94 2.99 2.93 2.86 2.79 1.64
Specifically, we also apply SVIP to EAGLE-2 (Li et al., 2024a), the current state-of-the-art (SOTA) speculative decoding system which utilizes the target model’s language modeling head on top of the draft model’s features to predict the next draft token, and dynamically constructs a draft tree at each draft position. Following Li et al. (2024a), we use Vicuna 7B, 13B (Chiang et al., 2023) as the base models, and set the sampling temperature to 1. To the best of our knowledge, we are also the first to investigate EAGLE-2’s effectiveness in long-form generation. The results of EAGLE-2 on MT-Bench are given in Table 2. Even on top of this SOTA speculative decoding system, SVIP yields consistent improvement, which is especially notable at longer context length, surpassing the vanilla EAGLE-2 by 14% speedup for Vicuna 7B and 7% for Vicuna 13B.
Table 3: Speedup of QwQ on MATH, AIME, and GPQA, along with their average generation length.
MATH500 GPQA AIME Avg
Level1 Level2 Level3 Level4 Level5
Avg. Length 1.3K 1.3K 1.6K 2.5K 3.6K 3.9K 6.2K
Const. 1.45 1.50 1.52 1.56 1.56 1.25 1.58 1.49
Heuristics 1.29 1.26 1.27 1.30 1.33 1.18 1.34 1.28
SVIP 1.65 1.68 1.75 1.78 1.82 1.52 1.77 1.71
3.2 Long-form Reasoning
Settings
Recently, o1-style reasoning models have come into the spotlight of LLM research. Thus, we are especially interested in seeing the effectiveness of SVIP and other speculative decoding strategies on such models, which often have very long outputs. Consequently, we utilize QwQ-32B-Preview (Team, 2024), the only applicable open-source reasoning model at the time of writing, which does not have off-the-shelf smaller variants, so we train our own draft model based on Qwen2.5 1.5B by distilling QwQ 32B on 1M mathematical Persona data (Chan et al., 2024) [^1] We have made this model publicly available and will provide links in camera-ready version.. Using this draft model, we conduct experiments on MATH (Hendrycks et al., 2021), AIME [^2] https://maa.org/maa-invitational-competitions/, and GPQA (Rein et al., 2023). We sample 200 questions from MATH ranging from level 1 to level 5, and use 73 questions released from 2022 to 2024 for AIME. For GPQA, we use the diamond test set.
SVIP achieves strong speedup in long-form reasoning
The overall results are given in Table 3. SVIP outperforms the two baselines by a large margin across different benchmarks and context lengths. Detailed analysis of the different length policies’ behaviours suggests similar results to the previous experiment: SVIP has an average proposal length similar to the constant draft length policy, but with a much higher draft token accept rate, leading to more effective speedup.
Figure 8: A case comparison of different length policies during one generation example. Drafts proposed by SVIP vary drastically in length, while constant and heuristics policies are insufficient to model such changes.
Table 4: The average draft model entropy and draft token acceptance rate of some representative tokens between QwQ-32B and 1.5B.
Token Avg. Entropy Accept Rate
All tokens 0.38 0.68
“Wait” 1.17 0.53
“Alright” 1.38 0.22
“Actually” 1.52 0.33
“)]” 0.12 1.00
“}}” 0.02 1.00
“ynomials” 0.01 1.00
“ponents” 0.01 1.00
In Table 4, we present some token cases with very low or very high acceptance rate. We find that completing subword units or equations is quite easy for the small draft model, while several keywords in QwQ’s reasoning patterns are much harder. The reverse correlation of draft entropy and acceptance rate on these tokens further validates the motivation of SVIP. In Figure 8, we plot the behaviors of different draft length policies in an example generation case. The drastic draft length oscillations in SVIP also highlight the importance of dynamic draft length policy.
@@ 4 Related Work
Unlike previous methods, SVIP  has no requirement for draft models except for autoregression, and is a totally training-free adaptive-length policy, which could boost any draft model’s performance.
@@
In contrast to previous methods which introduce tree expansion for proposing fixed fine-grained n-gram draft, our method SVIP involves no fixed tree expansion rules, and is mainly about a more dynamic and flexible draft length policy.
@@
Works in this category are few, but most relevant to ours. Liu et al. (2024) introduce PEARL, which lets the target model perform verification in parallel to draft generation, stopping the draft process when a mismatch is found. Huang et al. (2024) propose SpecDec++, which trains an acceptance prediction head on top of the draft model to predict the acceptance probability of the current draft token, stopping the draft round when the predicted acceptance probability falls below a constant threshold. Brown et al. (2024) propose Dynamic Depth Decoding (DDD) on top of EAGLE-2, which uses the sum of all tokens’ confidences in one level of its draft tree as an indicator to predict whether or not to continue draft generation. Concurrent with our work, Zhang et al. (2024) propose AdaEAGLE, which utilizes an MLP on top of EAGLE to predict the next round’s draft length.
In contrast to prior length-control strategies that necessitate the training of a length-prediction module, SVIP stands out with its training-free nature. This unique characteristic endows it with remarkable flexibility, allowing it to be seamlessly integrated and applied to any autoregressive draft model.
@@ 5 Conclusion
We propose SVIP, a flexible, training-free, and plug-and-play dynamic draft length policy for speculative decoding systems. Based on a theoretical lower bound of acceptance probability and its empirical approximation, SVIP determines whether to continue draft generation or to quit drafting based on the draft model’s entropy after the generation of each draft token. With extensive experiments spanning various base models, draft methods, test domains, and generation length, we validated the effectiveness of SVIP, sparking new insights on speculative decoding and more efficient large language models. For future work, we aim to investigate tighter bounds on the acceptance rate to improve the accuracy of acceptance probability estimates, thereby enabling more efficient draft length adaptation. Additionally, our current analysis may not fully capture context-dependent patterns. A more nuanced investigation into these patterns could further enhance performance.
Limitations
While SVIP advances speculative decoding through adaptive draft length control, it has several limitations that offer avenues for future work. The acceptance rate bound in Equation˜6 could be overly conservative. Developing a tighter bound would enhance the accuracy of acceptance probability estimates, enabling more effective draft length adaptation. Further, our analysis assumes they follow a simplified distribution discrepancy of SD systems. This may not fully capture the nuanced factors contributing to their occurrence, such as context-dependent patterns or model-specific biases. Context-dependent length proxy for SD systems could be the potential research direction.
@@ Appendix A The Complete Speculative Decoding Algorithms
[…]
We note that from Algorithm 2, it’s straightforward that the acceptance rate of a draft token $x_{t}$ in the sampling scenario is by definition $\min(1,\frac{p(x_{t})}{q(x_{t})})$.
@@ Appendix B Alternatives for Acceptance Rate Lower Bound Computation
$\geqslant 1-\sqrt{\frac{1}{2}\mathbb{KL}(q||p)}.$ (11)
@@
$\beta$ $\geqslant 1-\sqrt{1-e^{-\mathbb{KL}(q||p)}}.$ (12)
Compared with the Pinsker’s bound, it’s trivial to see that this bound is guaranteed to be always larger than 0. However, in practice we find that the Pinsker’s bound is about 11% tighter.
Appendix C $\gamma$ Approximation
C.1 Approximation Bound
Following Eq. (4), the acceptance rate $\beta$ satisfies:
$\beta\geq 1-\sqrt{\frac{1}{2}\mathbb{KL}(q||p)}=1-\sqrt{\frac{1}{2}H_{q,p}-\frac{1}{2}H_{q}}$
We denote the above bound as the actual bound. While this bound is theoretically sound, it relies on the exact access to the $H_{q,p}$ the cross entropy between target and the draft models, which is inaccessible during the SD drafting phase. To address this, approximate $H_{q,p}$ with $\gamma H_{q}$, where $\gamma$ is a random variable to describe the ratio between $H_{q,p}$ and $H_{q}$, i.e. $\gamma=H_{q,p}/H_{q}$, we could rewrite the bound:
$\mathbb{KL}(q||p)=H_{q,p}-H_{q}=(\gamma-1)H_{q}$
$\beta\geq 1-\sqrt{\frac{1}{2}(\gamma-1)H_{q}}$
To make this bound more practical, we approximate it using a constant $c$, obtaining
$\beta\geq 1-\sqrt{\frac{1}{2}(\gamma-1)H_{q}}\approx 1-\sqrt{cH_{q}}$
We denote the above bound as the approximation bound. Since $\gamma$ is a random variable, the tightness and reliability of this approximation depend on how well $c$ aligns with $\gamma$’s behavior. Specifically, we need our approximation bound is smaller than the actual bound:
$\beta\geq 1-\sqrt{\frac{1}{2}(\gamma-1)H_{q}}\geq 1-\sqrt{cH_{q}}$
Simplify the right side inequality:
$\gamma\leq 2c+1$ (13)
C.2 Theoretical Analysis
Now, from the $\gamma$’s distribution in Figure 5, let’s analyze the probability that the lower bounds hold by modeling $\gamma$’s distribution.
Gaussian Distribution
Let’s assume $\gamma\sim N(\mu,\sigma^{2})$, and the probability that the bound holds is:
$P(\gamma\leq 2c+1)=\Phi(\frac{2c+1-\mu}{\sigma})$ (14)
where $\Phi$ is the standard normal CDF. It demonstrates that:
  • If $c$ is small (e.g. $2c+1\leq\mu$), the probability that the bound holds is low.
  • If $c$ is large (e.g. $2c+1\geq\mu$), the bound holds with high probability; however the bound itself becomes loose.
Gamma Distribution
Given the right-skewed nature of $\gamma\geq 1$, we model as a shifted Gamma distribution: $\gamma=1+X$, where $X\sim\text{Gamma}(\alpha,\beta)$. The conditions for the bound to hold is :
$P(\gamma\leq 2c+1)=P(X\leq 2c)=\frac{\gamma(\alpha,\beta\cdot 2c)}{\Gamma(\alpha)}$ (15)
where $\gamma(\alpha,z)=\int_{0}^{z}t^{\alpha-1}e^{-t}dt$ is the lower incomplete gamma function, and $\Gamma(\alpha)$ is the gamma function. This probability depends on $c$, $\alpha$, and $\beta$.
  • If $c$ is small (e.g. $2c<\mathbb{E}[x]=\alpha/\beta$), so the probability that our approximation further lowers the actual bound is low.
  • It $c$ is large, the probability approaches 1, but the bound $\beta\geq 1-\sqrt{cH_{q}}$ is looser.
@@ Appendix D Additional Results on Long-form Generation
In Table 5, we present the results of greedy decoding using Qwen2.5 14B as target and 0.5B as draft, and find that the speedup ratio of greedy decoding is much higher compared with the sampling experiments in Section 3.1. Further investigation suggests that this is a result of repetition hallucination in both target and draft models during long-form greedy generation.
Table 5: Results of greedy decoding on MT-Bench. Greedy decoding leads to repetition hallucinations in both target and draft models in long-form generation, resulting in exaggerated speedup ratio.
Model Methods Context Length
@@
Qwen2.5 (14B, 0.5B) Const. 1.05 1.08 1.15 1.29 1.44 1.54 1.60 1.67